Jordan Company incurs annual fixed costs of $89,515. Variable costs for Jordan’s product are $21.35 per unit, and the sales price is $35.00 per unit. Jordan desires to earn an annual profit of $62,000.


Required
Use the contribution margin ratio approach to determine the sales volume in dollars and units required to earn the desired profit. (Do not round intermediate calculations. Round your final answers to the nearest whole number.)
Answer #1
Contribution Margin Ratio= (Sales - Variable cost )/ Sales
=(35-21.35)/35= 39%
Computation required sales for desired Profit
Required Contribution= Fixed Cost+ Desired Profit
( 89515+62000)/39%= $388500
Sales Volume in Unit= 388500/ $35= 11100 Unit
Sales Voulme in $= $388500